thrice a certain number added to 8 times the same number
Found 2 solutions by hyh_123, rapaljer:
Answer by hyh_123(8)

dont worry this a very easy question.
Thrice a certain number (x) means x*x*x=x^3(x to power 3)
Eight times the number is 8*x=8x
your equation would be
x^3+8x
=x(x^2+8)

Answer by rapaljer(4671)

In MY language, THRICE means THREE TIMES.

So, thrice a number added to 8 times a number means
3x + 8x
11x

Final answer!!